OpenStudy (anonymous):

Rewrite 3pi/32 in degree measure. Round to three decimal places.

OpenStudy (anonymous):

\[\frac{3 \pi }{32}=\frac{3*180}{32}=\frac{135}{8}=16.875 \text{ degrees}\]

OpenStudy (anonymous):

THANKS, :) sorry, to ask but why 180 (degrees?)? Is that because it 3pi is 2 revolutions plus a half revolution?

OpenStudy (saifoo.khan):

2pi is 360* 3/2 pi means 1.5 pi. Meaning 270*

OpenStudy (saifoo.khan):

And 1pi means 180*

OpenStudy (anonymous):

Thanks! Great! Makes sense! :)

OpenStudy (saifoo.khan):

No problem.

OpenStudy (anonymous):

1 each Pi is 180 degrees. 3*Pi is 3*180 degrees